# Regional and partial observability and control of waves

### v1

@@ Abstract.
Roughly speaking, this geometric condition requires that all rays of geometric optics in $\overline{\Omega}$, emanating from $\overline{\mathscr{O}}$, must reach the observation region $(0,T)\times\omega$. Our result significantly generalizes classical observability results, which recover the total energy of all solutions when the observation set $\omega$ satisfies the so-called Geometric Control Condition (GCC) a particular case corresponding to $\mathscr{O}=\Omega$.
@@
The proof of this previously unnoticed result combines a high-frequency observability estimate based on the propagation of singularities with a compactness-uniqueness argument that exploits the unique continuation properties of elliptic operators.
By duality, this observability result leads to new controllability results for the wave equation, ensuring that the projection of the solution onto $\mathscr{O}$ can be controlled by means of controls supported in $\omega$, with optimal spatial support.
@@ 1.1. Problem formulation
Let $\Omega$ be a bounded domain of ${\mathbb{R}}^{n}$, with boundary $\partial\Omega$ of class $\mathcal{C}^{\infty}$. We set
=×$\mathcal{L}={\mathbb{R}}\times\Omega\quad\text{and }\quad\partial\mathcal{L}={\mathbb{R}}\times\partial\Omega.$
We also introduce $A=(a_{ij}(x))$, a $n\times n$ matrix of $\mathcal{C}^{\infty}$ coefficients, symmetric, uniformly definite positive on a neighborhood of $\overline{\Omega}$, and we denote $\Delta_{A}=\sum_{i,j=1}^{n}\partial_{x_{j}}(a_{ij}(x)\partial_{x_{i}}\cdot)$ the corresponding Laplacian operator.
@@
(1.1) $\left\{\begin{array}[]{ll}\partial_{t}^{2}u-\Delta_{A}u=0,&\text{in }\,\,\mathcal{L},\\ u(t,.)=0,&\text{on }\,\,\partial\mathcal{L},\\ (u(0,\cdot),\partial_{t}u(0,\cdot))=(u_{0},u_{1}).&\end{array}\right.$
@@ 1.2. Main results
Given the domain $\Omega$, the observation subdomain $\omega\subset\Omega$, and the time-horizon $T>0$, let the subdomain $\mathscr{O}$ be such that every geodesic ray emanating from $\overline{\mathscr{O}}$ intersects $\omega$ before the time $T$.
@@
This definitive result holds for all subdomains $\omega$ and initial data with support in $\overline{\mathscr{O}}$, under the condition that the pair $(\omega,\mathscr{O})$ fulfills a mutual or joint microlocal relation, so that all rays emanating from $\overline{\mathscr{O}}$ reach $(0,T)\times\omega$. It is a natural extension of the classical one guaranteeing the observability of all solutions, namely the GCC, which corresponds to the very particular case $\mathscr{O}=\Omega$ in our setting. Indeed, it is natural that, when the support of the initial data lies in $\overline{\mathscr{O}}$, its observation only depends on the dynamics of the rays emanating from $\overline{\mathscr{O}}$, independently of what other rays (the ones departing away from $\overline{\mathscr{O}}$) do.
@@
These results enter in the context of “enlarged observability/controllability” introduced in [12], according to which when limiting the class of solutions under consideration the requirements for observability can be weakened. However, in [12], because of the use of the multiplier method, improvements were only achieved at the level of the needed observability time. The results of the present paper constitute a much more precise answer to these questions.
A similar result holds when the observation is done along the boundary. For this, we need the notion of nondiffractive points of the boundary that will be detailed in Definition 3.1, see also [1, Definition, pp.1037] .
@@ 1.3. Examples
A $1$-d example. The microlocal assumption on the pair $(\omega,\mathscr{O})$ in Theorem 1.1, is, in general, much weaker than the GCC since it only concerns the rays emanating from $\mathscr{O}$. It also requires a shorter observation time. This is even the case in $1$-d.
@@
By symmetry considerations, it is then easy to check that Theorem 1.1 holds for any $T>1$. This minimal time corresponds to the arrival in $\omega$ of a characteristic starting from $x=-1/4$, propagating to the right.
However, the classical, sharp condition for unique continuation in (1.9) or the GCC require $T>T_{UC}=3/2$. This condition is indeed optimal when aiming to observe all solutions since one can build initial data localized in $(-3/4,-3/4+\varepsilon)$ ($\varepsilon>0$ small) leading to waves propagating towards the right at speed one, and vanishing in $\omega$ during the time interval $(0,3/2-\varepsilon)$.
Therefore, the global observability estimates (1.4) or (1.6) do not hold for all initial data for the intermediate times $21<T<3/$ 2, but, according to our result, they do hold for initial data localized in $\overline{\mathscr{O}}$.
@@
    Another example, illustrated in Figure 3, still when $\Omega$ is the unit ball of ${\mathbb{R}}^{2}$, corresponds to an observation subdomain $\omega$ which is an $\varepsilon$ neighborhood of one third of the boundary of $\Omega$, with angles in $(-\pi/3,\pi/3)$, and $\mathscr{O}$ located in the vertical strip $\{x=(x_{1},x_{2})\in\Omega\hbox{ with }x_{1}\leqslant\cos(\alpha)\}$ for some $\alpha\in(\pi,4\pi/3)$. The critical time $T$ given by Theorem 1.1 is finite but GCC fails, whatever the time-horizon is, due to the vertical diameter, corresponding to a geodesic ray which bounces back and forth endless, without ever entering the observation set $\omega$.
@@ 1.4. Some relevant consequences
Note that, if we further assume that the time horizon $T$ satisfies (1.9), we can find controls that simultaneously assure the exact control of the projection over $\mathscr{O}$ and the approximate controllability property everywhere in the domain $\Omega$, see Remark 5.3.
The problem of controllability in the absence of GCC has been the object of extensive study, see, for instance, [22, 10, 20, 9]. In those papers, one aims to quantify the property of approximate controllability, by identifying subspaces of initial data that can actually be controlled. These spaces are typically very small, imposing suitable analyticity restrictions. Surprisingly, our paper is the first one in which, in the absence of GCC, the microlocal properties of the pair $(\omega,\mathscr{O})$ are exploited to identify a class of controllable initial data in the sharp energy spaces.
This definitive result, valid for arbitrary observation sets $\omega$, generalizes and encompasses as a particular case the classical setting in which $\mathscr{O}=\Omega$, that is, when observability is required for all solutions of the wave equation. In this case, our condition naturally reduces to the well-known Geometric Control Condition (GCC).
@@
By duality, this leads to novel control results for the wave equation, ensuring the control of the projections over $\mathscr{O}$, or its sharper microlocal counterpart, by means of controls localized in $\omega$, up to some smoother error terms (see Section 5 for the precise statement).
Throughout the paper, we also present several illustrative examples involving different domains $\Omega$, observation subsets $\omega$, and time horizons $T$, which do not satisfy the classical GCC. Nonetheless, our result yields new observability estimates and control results in these cases, highlighting its broader applicability beyond the classical framework.
@@
As an interesting corollary of our novel observability result, we derive new unique continuation properties for specific classes of solutions — such as those with initial data supported in $\mathscr{O}$ — even in settings where the classical condition (1.9) fails. Consequently, these results apply in situations where existing uniqueness theorems of Holmgren do not suffice.
@@ 2. Proofs of the main results
We essentially focus on the proof of Theorem 1.1, i.e.,of the estimate (1.6), since the proof of estimate (1.4) is similar, and is thus left to the reader, with the indication of the additional steps needed. The general strategy of the proof follows the program described in Section 1.5.
@@
Under the assumptions of Theorem 1.1, consider a solution $u$ to system (1.1) with initial data $(u_{0},u_{1})\in H^{-1}(\Omega)\times(H^{2}\cap H^{1}_{0})^{\prime}(\Omega)$ supported in $\overline{\mathscr{O}}$, and satisfying $u\in L^{2}(\omega_{T})$. Then $(u_{0},u_{1})\in L^{2}(\Omega)\times H^{-1}(\Omega)$ and $u\in\mathcal{C}^{0}({\mathbb{R}},L^{2}(\Omega))\cap\mathcal{C}^{1}({\mathbb{R}},H^{-1}(\Omega))$.
@@
We then deduce the following corollary, which fulfills the first step of the proof, providing a first observability estimate with a compact reminder.
@@
The proof of estimate (2.1) is similar, except that it relies on the propagation of the $L^{2}$-wave front set along the bicharacteristic flow. ∎
@@
At this point it is worth noticing that this uniqueness result is not standard since it only applies to the solutions with initial data of support in $\overline{\mathscr{O}}$. It is not a consequence of Holmgren’s uniqueness theorem nor any of its generalisations, but rather a corollary of the Lemma 2.2, which establishes a relaxed version of the observability inequalities we aim, with an added compact additive remainder term.
@@
Similarly as in the proof of Lemma 2.2, we focus on finite energy solutions, the case of weaker solutions being similar.
@@
Moreover, thanks to the estimate (2.2) in Lemma 2.2, it is clear that $\mathcal{N}\subset H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ and
@@
Using then the compact embedding $H^{1}_{0}(\Omega)\hookrightarrow L^{2}(\Omega)$  [^1] In fact, since we are considering data which are supported in $\overline{\mathscr{O}}$, we can also consider the compact embedding of $H^{1}_{0}(\mathscr{O})\hookrightarrow L^{2}(\mathscr{O})$. This remark allows to deduce the same results even in cases in which $\Omega$ is unbounded, provided $\mathscr{O}$ is bounded. we deduce that $\mathcal{N}(T)$ has a finite dimension. In addition, the matrix operator ${\mathcal{A}}=\begin{pmatrix}0&1\\ \Delta_{A}&0\end{pmatrix}$ defines a linear bounded operator in $\mathcal{N}$. This is so since the wave equation, written on the column vector unknown $U=\begin{pmatrix}u\\ u_{t}\end{pmatrix}$ takes the form $U_{t}={\mathcal{A}}U$. Therefore, ${\mathcal{A}}$ operates continuously in $\mathcal{N}$ and corresponds to the application of the time derivative on the solutions of the wave equation (1.1), transferring the initial data $(u_{0},u_{1})$ into $(u_{1},\Delta_{A}u_{0}).$
If $\mathcal{N}$ were non-trivial, the operator ${\mathcal{A}}$ would have an eigenvalue. But, as we shall see, this is impossible, concluding, by contradiction, that $\mathcal{N}=\{(0,0)\}$.
Indeed, if $\lambda\in{\mathbb{C}}$ is an eigenvalue and $(u_{0},u_{1})\neq(0,0)$ in $H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ is an eigenvector, we have,
$\Delta_{A}u_{0}-\lambda^{2}u_{0}=0\text{ in }\Omega,\qquad u_{0}=0\text{ on }\partial\Omega,\qquad\hbox{ and }\qquad\lambda u_{0}=0\text{ in }\omega.$
@@
We are now in conditions to conclude the proof of Theorem 1.1, i. e. of estimate (1.6).
@@
The sequence $(u_{k})$ is bounded in the energy space $C^{0}([0,T];H^{1}_{0}(\Omega))\cap C^{1}([0,T];L^{2}(\Omega))$. Thus, after extracting a subsequence, we may assume that it converges weakly in $H^{1}(\mathcal{L}_{T})$ to another solution $u\in H^{1}(\mathcal{L}_{T})$ of (1.1), corresponding to an initial datum $(u_{0},u_{1})\in H^{1}_{0}(\Omega)\times L^{2}(\Omega)$ with support in $\overline{\mathscr{O}}$, weak limit of $(u_{0,k},u_{1,k})$ in the energy space.
@@
But the right hand-side tends to 00 as $k\to\infty$ thanks to (2.3) and the fact that $(u_{0,k},u_{1,k})\to(0,0)$ in $L^{2}(\Omega)\times H^{-1}(\Omega)$.  [^2] Note that, here again, the same argument applies if $\Omega$ is unbounded but $\mathscr{O}$ is bounded, since the compact embedding of $H^{1}_{0}(\mathscr{O})\times L^{2}(\mathscr{O})$ into $L^{2}(\mathscr{O})\times H^{-1}(\mathscr{O})$ can be employed. This yields a contradiction.

### v2

@@ Abstract.
Roughly speaking, this geometric condition requires that all rays of geometric optics in $\overline{\Omega}$, emanating from $\overline{\mathscr{O}}$, must reach the observation region $(0,T)\times\omega$. Our result generalizes classical observability results, which recover the total energy of all solutions when the observation set $\omega$ satisfies the so-called Geometric Control Condition (GCC)—a particular case corresponding to $\mathscr{O}=\Omega$.
@@
The proof of our result combines a high-frequency observability estimate—based on the propagation of singularities—with a compactness-uniqueness argument that exploits the unique continuation properties of elliptic operators.
By duality, this observability result leads to controllability results for the wave equation, ensuring that the projection of the solution onto $\mathscr{O}$ can be controlled by means of controls supported in $\omega$, with optimal spatial support.
@@ 1.1. Problem formulation
Let $\Omega$ be a bounded domain of ${\mathbb{R}}^{n}$, with boundary $\partial\Omega$ of class $\mathscr{C}^{\infty}$. We set
[…]
We also introduce $A=(a_{ij}(x))$, a $n\times n$ matrix of $\mathscr{C}^{\infty}$ coefficients, symmetric, uniformly definite positive on a neighborhood of $\overline{\Omega}$, and we denote $\Delta_{A}=\sum_{i,j=1}^{n}\partial_{x_{j}}(a_{ij}(x)\partial_{x_{i}}\cdot)$ the corresponding Laplacian operator.
@@
[…]
@@ 1.2. Main results
Given the domain $\Omega$, the observation subdomain $\omega\subset\Omega$, and the time-horizon $T>0$, let the subdomain $\mathscr{O}$ be such that every generalized bicharacteristic ray (see Definition 3.2 for their precise definition) starting from $\{t=0\}\times\overline{\mathscr{O}}$ intersects the set $(0,T)\times\omega$.
@@
This result holds for all subdomains $\omega$ and initial data with support in $\overline{\mathscr{O}}$, under the condition that the pair $(\omega,\mathscr{O})$ fulfills a mutual or joint microlocal relation, so that all rays emanating from $\overline{\mathscr{O}}$ reach $(0,T)\times\omega$. It is a natural extension of the classical one guaranteeing the observability of all solutions, namely the GCC, which corresponds to the case $\mathscr{O}=\Omega$ in our setting. Indeed, it is natural that, when the support of the initial data lies in $\overline{\mathscr{O}}$, its observation only depends on the dynamics of the rays emanating from $\overline{\mathscr{O}}$, independently of what other rays (the ones departing away from $\overline{\mathscr{O}}$) do.
@@
After this work has been submitted, the referee kindly pointed out to us the work [1, Section 2], which deals with a similar problem in a particular setting in which $\mathscr{O}$ is a ball $B(r)$ strictly included in $\Omega$ and the metric is flat, under a slightly more demanding condition. Namely, the article [1] requires that there is a neighborhood $B(r+\delta)$ ($\delta>0$) of the ball for which every generalized bicharacteristic ray passing through $0,\delta]\times B(r+\delta)$ intersects $(0,T)\times\omega$ (or $(0,T)\times\Gamma$ at a non-diffractive point if $\Gamma$ is a part of the boundary on which the observation is done, see Theorem 1.4 for the counterpart of Theorem 1.1 when the observation is performed on the boundary). Using partition of unity, we could derive from [1] cases in which $\mathscr{O}$ is strictly included in $\Omega$, but it will not be sufficient to handle cases in which $\overline{\mathscr{O}}$ touches the boundary $\partial\Omega$ as in the example presented in Figure 3 afterwards.
@@
These results enter in the context of “enlarged observability/controllability” introduced in [13], according to which when limiting the class of solutions under consideration the requirements for observability can be weakened. However, in [13], because of the use of the multiplier method, improvements were only achieved at the level of the needed observability time. The results of the present paper constitute a more precise answer to these questions.
A similar result holds when the observation is done along the boundary. For this, we need the notion of nondiffractive points of the boundary that will be detailed in Definition 3.1, see also [2, Definition, pp.1037] .
@@ 1.3. Examples
A $1$-d example. The microlocal assumption on the pair $(\omega,\mathscr{O})$ in Theorem 1.1, is, in general, weaker than the GCC since it only concerns the rays emanating from $\overline{\mathscr{O}}$. It also requires a shorter observation time. This is even the case in $1$-d.
@@
By symmetry considerations, it is then easy to check that Theorem 1.1 holds for any $T>1$. This minimal time corresponds to the arrival in $\omega$ of a characteristic starting from x=1/4x=-1/4, propagating to the right.
However, the classical, sharp condition for unique continuation in (1.9) (in the following, $T_{UC}$ stands for the critical time for unique continuation; it is given by $2\sup_{x\in\Omega}d(x,\omega)$ from the Hörmander-Tataru-Robbiano-Zuily uniqueness theorem [25, 22, 8]) or the GCC require $T>T_{UC}=3/2$ . This condition is indeed optimal when aiming to observe all solutions since one can build initial data localized in $(-3/4,-3/4+\varepsilon)$ ($\varepsilon>0$ small) leading to waves propagating towards the right at speed one, and vanishing in $\omega$ during the time interval $(0,3/2-\varepsilon)$.
Therefore, the global observability estimates (1.4) or (1.6) do not hold for all initial data for the intermediate times $1<T<3/2$, but, according to our result, they do hold for initial data localized in $\overline{\mathscr{O}}$.
@@
    Another example, illustrated in Figure 3, still when $\Omega$ is the unit ball of ${\mathbb{R}}^{2}$, corresponds to an observation subdomain $\omega$ which is an $\varepsilon$ neighborhood of one third of the boundary of $\Omega$, with angles in $(-\pi/3,\pi/3)$, and $\mathscr{O}$ being the vertical strip $\{x=(x_{1},x_{2})\in\Omega\hbox{ with }x_{1}<\cos(\alpha)\}$ for some $\alpha\in(\pi,4\pi/3)$. The critical time $T$ given by Theorem 1.1 is finite but GCC fails, whatever the time-horizon is, due to the vertical diameter, corresponding to a geodesic ray which bounces back and forth endless, without ever entering the observation set $\omega$.
@@ 1.4. Some relevant consequences
Note that, if we further assume that the time horizon $T$ is sufficiently large to guarantee the injectivity of the operator $\mathrm{Obs}_{T}$ in (1.2) (in fact, this corresponds to the condition $T>2\sup d(x,\omega)$, see (1.9) afterwards), we can find controls that simultaneously assure the exact control of the projection over $\mathscr{O}$ and the approximate controllability property everywhere in the domain $\Omega$, see Remark 5.3.
The problem of controllability in the absence of GCC has been the object of extensive study, see, for instance, [23, 11, 21, 10]. In those papers, one aims to quantify the property of approximate controllability, by identifying subspaces of initial data that can actually be controlled. These spaces are typically very small, imposing suitable analyticity restrictions. Surprisingly, analyzing what can be controlled for solutions of the wave equation in the usual energy space in the absence of GCC has not received much attention, and, besides our work, we are only aware of the work [1] focusing on a very particular case.
Our result, valid for arbitrary observation sets $\omega$, encompasses as a particular case the classical setting in which $\mathscr{O}=\Omega$, that is, when observability is required for all solutions of the wave equation. In this case, our condition naturally reduces to the well-known Geometric Control Condition (GCC).
@@
By duality, these observation type properties yield control results for the wave equation in which the control objective is to control the projection of the solutions over $\mathscr{O}$, or its microlocal projection up to some smoother error terms (see Section 5 for the precise statement).
Throughout the paper, we also present several illustrative examples involving different domains $\Omega$, observation subsets $\omega$, and time horizons $T$, which do not satisfy the classical GCC. Nonetheless, our results apply in those cases, as our goal is to discuss all the information on the solutions of the wave equation that can be derived from a given observation set $(0,T)\times\omega$.
@@
As an interesting corollary of our observability result, we derive unique continuation properties for specific classes of solutions — such as those with initial data supported in $\mathscr{O}$ — even in settings where the classical condition (1.9) fails. Consequently, these results apply in situations where existing uniqueness theorems of Holmgren do not suffice.
@@ 1.5. Methodology of proof
Acknowledgments. The authors thank the referee for having pointed out the article [1].
@@ 2. Proofs of the main results
We essentially focus on the proof of Theorem 1.1, i.e., of the estimate (1.6), since the proof of estimate (1.4) is similar, and is thus left to the reader, with the indication of the additional steps needed, see Section 2.2 afterwards. The general strategy of the proof follows the program described in Section 1.5.
@@
Under the assumptions of Theorem 1.1, consider a solution $u$ to system (1.1) with initial data $(u_{0},u_{1})\in H^{-1}(\Omega)\times(H^{2}\cap H^{1}_{0})^{\prime}(\Omega)$ supported in $\overline{\mathscr{O}}$, and satisfying $u\in L^{2}(\omega_{T})$. Then $(u_{0},u_{1})\in L^{2}(\Omega)\times H^{-1}(\Omega)$ and $u\in\mathscr{C}^{0}({\mathbb{R}},L^{2}(\Omega))\cap\mathscr{C}^{1}({\mathbb{R}},H^{-1}(\Omega))$.
@@
We then deduce the following corollary, which fulfills the first step of the proof, providing a first observability estimate with a compact remainder.
@@
The proof of estimate (2.1) is similar, except that it relies on the propagation of the $L^{2}$-wave front set along the bicharacteristic flow, that is the first item of Lemma 2.1. ∎
@@
At this point it is worth noticing that this uniqueness result is not standard since it only applies to the solutions with initial data supported in $\overline{\mathscr{O}}$. It is not a consequence of Holmgren’s uniqueness theorem nor any of its generalisations, but rather a corollary of Corollary 2.2, which establishes a relaxed version of the observability inequalities we aim, with an added compact additive remainder term.
@@
Similarly as in the proof of Corollary 2.2, we focus on finite energy solutions, the case of weaker solutions being similar.
@@
Thanks to the estimate (2.2) in Corollary 2.2, it is clear that $\mathcal{N}\subset H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ and
@@
Using then the compact embedding $H^{1}_{0}(\Omega)\hookrightarrow L^{2}(\Omega)$  [^1] In fact, since we are considering data which are supported in $\overline{\mathscr{O}}$, we can also consider the compact embedding of $H^{1}_{0}(\mathscr{O})\hookrightarrow L^{2}(\mathscr{O})$. This remark allows to deduce the same results even in cases in which $\Omega$ is unbounded, provided $\mathscr{O}$ is bounded. we deduce that $\mathcal{N}$ has a finite dimension. In addition, the matrix operator ${\mathcal{A}}=\begin{pmatrix}0&1\\ \Delta_{A}&0\end{pmatrix}$ defines a linear bounded operator in $\mathcal{N}$. This is so since the wave equation, written on the column vector unknown $U=\begin{pmatrix}u\\ u_{t}\end{pmatrix}$ takes the form $U_{t}={\mathcal{A}}U$. Therefore, $\mathcal{A}$} operates continuously in $\mathcal{N}$ and corresponds to the application of the time derivative on the solutions of the wave equation (1.1), transferring the initial data $(u_{0},u_{1})$ into $(u_{1},\Delta_{A}u_{0}).$
If $\mathcal{N}$ were non-trivial, the operator $\mathcal{A}$} would have an eigenvalue corresponding to an eigenvector in $\mathcal{N}$. But, as we shall see, this is impossible, concluding, by contradiction, that $\mathcal{N}=\{(0,0)\}$.
Indeed, if $\lambda\in{\mathbb{C}}$ is an eigenvalue and $(u_{0},u_{1})\neq(0,0)$ in $H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ is an eigenvector of $\mathcal{A}$ in $\mathcal{N}$, we have,
[…]
@@
We are now in conditions to conclude the proof of Theorem 1.1, i.e., of estimate (1.6).
@@
The sequence $(u_{k})$ is bounded in the energy space $\mathscr{C}^{0}([0,T];H^{1}_{0}(\Omega))\cap\mathscr{C}^{1}([0,T];L^{2}(\Omega))$. Thus, after extracting a subsequence, we may assume that it converges weakly in $H^{1}(\mathcal{L}_{T})$ to another solution $u\in H^{1}(\mathcal{L}_{T})$ of (1.1), corresponding to an initial datum $(u_{0},u_{1})\in H^{1}_{0}(\Omega)\times L^{2}(\Omega)$ with support in $\overline{\mathscr{O}}$, weak limit of $(u_{0,k},u_{1,k})$ in the energy space.
@@
But the right hand-side tends to 0 as $k\to\infty$ thanks to (2.3) and the fact that $(u_{0,k},u_{1,k})\to(0,0)$ in $L^{2}(\Omega)\times H^{-1}(\Omega)$.  [^2] Note that, here again, the same argument applies if $\Omega$ is unbounded but $\mathscr{O}$ is bounded, since the compact embedding of $H^{1}_{0}(\mathscr{O})\times L^{2}(\mathscr{O})$ into $L^{2}(\mathscr{O})\times H^{-1}(\mathscr{O})$ can be employed. This yields a contradiction.